OpenStudy (anonymous):

the area of a trapezoid is A=0.5(b1+b2) h Find b1 if a= 100, h=5 and b2 =12

OpenStudy (anonymous):

b1+b2=A/0.5h b1+12=100/0.5*5 b1=(1000/25)-12 b1=

OpenStudy (anonymous):

28

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thank you!